OpenStudy (anonymous):

Find an equation of the line that is tangent to the graph of f and parallel to the given line. Function f(x) = x2 − x Line x + 3y − 6 = 0

OpenStudy (anonymous):

function \(f(x)=x^2-x\) had derivative \(f'(x)=2x-1\)

OpenStudy (anonymous):

slope of \(x+3y-6=0\) is \(-\frac{1}{3}\)

OpenStudy (anonymous):

set them equal and solve for \(x\)

OpenStudy (anonymous):

2x-1=-1/3